# Eigenvalue analysis of three-state quantum walks with general coin matrices

### v4

@@ III.3 Eigenvalue analysis for two-phase QW with defects
Here $\times$ denotes the cross product on $\mathbb{C}^{2}$ defined by $[x_{1}\ x_{2}]\times[y_{1}\ y_{2}]=x_{1}y_{2}-x_{2}y_{1}$. Additionally, the multiplicity of $e^{i\lambda}\in\sigma_{p}(U)$ is $1$ with a unique eigenvector $\iota^{-1}\tilde{\Psi}$ (up to global phase) given by:
$\tilde{\Psi}(x)=$$\begin{cases}(\zeta_{\infty}^{<})^{x-x_{+}}T_{[x_{-},x_{+}-1]}\ket{v_{-\infty}^{>}},&x_{+}\leq x,\\ T_{[x_{-},x-1]}\ket{v_{-\infty}^{>}},&x\in(x_{-},x_{+}),\\ (\zeta_{-\infty}^{>})^{x-x_{-}}\ket{v_{-\infty}^{>}},&x\leq x_{-}.\end{cases}$
@@ III.3.1 Homogeneous model
$C_{x}=\begin{bmatrix}a^{(1,1)}&a^{(1,2)}&a^{(1,3)}\\ a^{(2,1)}&a^{(2,2)}&a^{(2,3)}\\ a^{(3,1)}&a^{(3,2)}&a^{(3,3)}\end{bmatrix}.$
@@
$\sigma_{p}(U)=\begin{cases}\left\{\frac{\overline{a_{33}}}{a_{11}}e^{i\Delta}\right\}&\left(\frac{a_{33}}{\overline{a_{11}}}\right)^{2}=\frac{a_{12}a_{21}}{\overline{a_{32}}\overline{a_{23}}},\\ \emptyset&otherwise.\end{cases}$
@@
$\ \tilde{\Psi}(x)=k_{x}\begin{bmatrix}a^{(3,3)}\overline{a^{(3,2)}}\\ \overline{a^{(1,1)}}a^{(2,1)}\end{bmatrix}$
@@
From the main theorem, we can state that for $\lambda\in[0,2\pi)$, $e^{i\lambda}\neq\frac{\overline{a_{33}}}{a_{11}}e^{i\Delta}$, (case $A(\lambda)\neq 0$) is an eigenvalue of $U$ if and only if $|\mathrm{tr}(T(\lambda))|>2$ and there exists $\tilde{\Psi}:\mathbb{Z}\rightarrow\mathbb{C}^{2}\setminus\{\mathbf{0}\}$ such that
@@
for $\tilde{\Psi}$ to be exponentially decaying on both the left and the right. However, since $\zeta^{<}\neq\zeta^{>},$ $\tilde{\Psi}(0)=\mathbf{0}$ and this yields $\tilde{\Psi}=\mathbf{0}$.  Thus only $\frac{\overline{a_{33}}}{a_{11}}e^{i\Delta}$ (case $A(\lambda)=0$) could be an eigenvalue.  From the main theorem, $e^{i\lambda}=\frac{\overline{a_{33}}}{a_{11}}e^{i\Delta}$ is an eigenvalue if and only if $\left(\frac{a^{(3,3)}}{\overline{a^{(1,1)}}}\right)^{2}=\frac{a^{(1,2)}a^{(2,1)}}{\overline{a^{(3,2)}}\overline{a^{(2,3)}}}$ and the corresponding eigenvector becomes  $\iota^{-1}\tilde{\Psi}$ such that for each $x\in\mathbb{Z}$, there exists $k_{x}\in\mathbb{Z}$ such that
@@
This result confirms that localization does not generally occur in homogeneous three-state QWs unless specific algebraic conditions are satisfied. Even when an eigenvalue exists, the corresponding eigenstates are fundamentally different from those obtained via the transfer matrix method in the inhomogeneous case: they are not exponentially decaying but instead are compactly supported (nonzero only at finitely many positions). This observation aligns with earlier findings in the literature, such as those derived via Fourier analysis in [29], and further motivates the study of inhomogeneous models, where localization, characterized by exponentially decaying eigenstates, emerges more generically.
@@ III.3.2 One-defect model
The one-defect model introduces a localized perturbation to an otherwise homogeneous QW. Specifically, the coin matrix is modified at a single site, typically taken to be the origin $x=0$, while the coin is kept uniform elsewhere. This minimal breaking of translational invariance is sufficient to drastically alter the spectral and dynamical properties of the system. The one-defect model is particularly important in the study of QWs because it serves as a simple yet rich example of how inhomogeneities lead to localization and bound states. The one-defect model is defined such that $C_{+\infty}=C_{-\infty}=C$, and
$C_{x}=\begin{cases}C_{0}&x=0,\\ C&x\neq 0,\end{cases}$
where we denote each element of $C$ as
$C=\begin{bmatrix}a^{(1,1)}&a^{(1,2)}&a^{(1,3)}\\ a^{(2,1)}&a^{(2,2)}&a^{(2,3)}\\ a^{(3,1)}&a^{(3,2)}&a^{(3,3)}\end{bmatrix},\ e^{i\Delta}=\det C.$
@@
$\ \tilde{\Psi}(0)=k_{1}\begin{bmatrix}\overline{a^{(1,1)}}a^{(1,2)}\\ a^{(3,3)}\overline{a^{(2,3)}}\end{bmatrix},\ \tilde{\Psi}(1)=k_{2}\begin{bmatrix}a^{(3,3)}\overline{a^{(3,2)}}\\ \overline{a^{(1,1)}}a^{(2,1)}\end{bmatrix}$
@@
Figure 2: Probability distribution $\mu_{t}^{(\Psi_{0})}(x)$ for the one-defect Fourier walk with different values of $\theta$ at time $100$ with initial state $\Psi_{0}(0)=[\frac{1}{\sqrt{3}},\frac{i}{\sqrt{3}},\frac{1}{\sqrt{3}}]^{t}$ and $\Psi_{0}(x)=\mathbf{0}$ for $x\neq 0$. All of (a-d) have a peak around the origin, which indicates localization. The result shown in Fig. 1 corresponds to these localization.
@@ III.3.3 Two-phase model
Lastly, we consider the two-phase model, where the time evolution differs between the left and right halves of the lattice. That is, the coin matrix takes two distinct values depending on whether $x<0$ or $x\geqslant 0$. This model has been widely studied in the context of quantum localization and topological insulators, where phase boundaries can give rise to localized edge states. The two-phase model is defined such that $C_{+\infty}=C_{-\infty}=C$, and
$C_{x}=\begin{cases}C_{+\infty}&x\geqslant 0,\\ C_{-\infty}&x<0.\end{cases}$
@@
For $\lambda\in[0,2\pi)$, $e^{i\lambda}$ is an eigenvalue of $U$ if and only if there exists $\tilde{\Psi}:\mathbb{Z}\rightarrow\mathbb{C}^{2}$ such that Conditions 1, 2, and 3 are all satisfied.
@@
$\ \tilde{\Psi}(0)=k\begin{bmatrix}\overline{a_{-\infty}^{(1,1)}}a_{-\infty}^{(1,2)}\\ a_{-\infty}^{(3,3)}\overline{a_{-\infty}^{(2,3)}}\end{bmatrix},$
@@
$\ \tilde{\Psi}(x)=\begin{cases}k_{x}\begin{bmatrix}a_{-\infty}^{(3,3)}\overline{a_{-\infty}^{(3,2)}}\\ \overline{a_{-\infty}^{(1,1)}}a_{-\infty}^{(2,1)}\end{bmatrix}&\left(\frac{a_{-\infty}^{(3,3)}}{\overline{a_{-\infty}^{(1,1)}}}\right)^{2}=\frac{a_{-\infty}^{(1,2)}a_{-\infty}^{(2,1)}}{\overline{a_{-\infty}^{(3,2)}}\overline{a_{-\infty}^{(2,3)}}},\\ \mathbf{0}&otherwise.\end{cases}$
@@ IV Numerical analysis on Fourier walk
Figure 4: Probability distribution $\mu_{t}^{(\Psi_{0})}(x)$ for the two-phase Fourier walk with different values of $\theta$ at time $100$ with initial state $\Psi_{0}(0)=[\frac{1}{\sqrt{3}},\frac{i}{\sqrt{3}},\frac{1}{\sqrt{3}}]^{t}$ and $\Psi_{0}(x)=\mathbf{0}$ for $x\neq 0$. All of (a-d) have a peak around the origin, which indicates localization. The result shown in Fig. 3 corresponds to these localization.

### v5

@@ III.1 Eigenvalue and Eigenvector
Here $\times$ denotes the cross product on $\mathbb{C}^{2}$ defined by $x_{1}\ x_{2}]\times[y_{1}\ y_{2}]=x_{1}y_{2}-x_{2}y_{1}$. Additionally, the multiplicity of $e^{i\lambda}\in\sigma_{p}(U)$ is $1$ with a unique eigenvector $\iota^{-1}\tilde{\Psi}$ (up to global phase) given by:
[…]
As a remark, we note that the cross product in two dimensions is equivalent to the determinant. Thus, the condition $\chi(\lambda)=0$ can be rewritten as:
[…]
III.2 Essential spectrum
We now describe the spectrum of the unitary operator $U$, denoted by $\sigma(U)$, for the two-phase three-state quantum walk with finitely many defects (including its continuous part). By standard spectral theory, the spectrum admits the disjoint decomposition
$\sigma(U)=\sigma_{\mathrm{disc}}(U)\ \dot{\cup}\ \sigma_{\mathrm{ess}}(U),$ (9)
[…]
The essential spectrum is unchanged by finite defects and is determined solely by the walk’s limiting behavior at $\pm\infty$. Thus, for our two-phase QW with finite defects model, the essential spectrum of the inhomogeneous walk $U$ can be expressed in terms of the tail operators:
$\sigma_{\mathrm{ess}}(U)=\sigma(U_{-\infty})\cup\sigma(U_{+\infty}).$ (10)
Here $U_{-\infty},\ U_{+\infty}$ are homogeneous walks with constant coins $\mathsf{C}_{-\infty}$ and $\mathsf{C}_{+\infty}$, respectively. For details, we refer to the general results in [49, 50].
Since the tail walks $U_{\pm\infty}$ are homogeneous (translation-invariant), their spectra admit the standard fiber decomposition under the Fourier transform [29], so the continuous part is purely absolutely continuous [49]. Moreover, any point spectrum of a homogeneous walk can only come from a $k$-independent (flat-band) eigenvalue of the fiber matrix, and hence has infinite multiplicity. Consequently,
$\sigma(U_{\pm\infty})=\sigma_{\mathrm{ac}}(U_{\pm\infty})\ \cup\ \sigma^{\infty}_{\mathrm{ess}}(U_{\pm\infty}),$ (11)
where $\sigma_{\mathrm{ac}}(U_{\pm\infty})$ is absolutely continuous and $\sigma^{\infty}_{\mathrm{ess}}(U_{\pm\infty})$ consists (if nonempty) of eigenvalues of infinite multiplicity. Our transfer-matrix formalism captures $\sigma^{\infty}_{\mathrm{ess}}(U_{\pm\infty})$ via Corollary 2, which explicitly derives the corresponding eigenvalues and shows that they arise only at the exceptional points $\Lambda_{0}$ defined in Eq. 8.
The absolutely continuous part can be derived using Weyl’s Criterion for the essential spectrum of $U_{\pm\infty}$ with non-exceptional points $\lambda\notin\Lambda_{0}$. A value $e^{i\lambda}$ belongs to the essential spectrum $\sigma_{\mathrm{ess}}(U)$ if and only if there exists a sequence of unit vectors $\{\Psi_{n}\}_{n=1}^{\infty}\subset\mathcal{H}$ (a Weyl sequence) such that:
  1. 1.
    $\|\Psi_{n}\|=1$ for all $n$,
  2. 2.
    $\Psi_{n}$ converges weakly to 0,
  3. 3.
    $\|(U-e^{i\lambda})\Psi_{n}\|\to 0$ as $n\to\infty$.
Equivalently, for $\tilde{\Psi}_{n}:=\iota\Psi_{n}\in\ell^{2}(\mathbb{Z};\mathbb{C}^{2})$, since $\iota$ is bijective, the condition is equivalent to the existence of a sequence $\{\tilde{\Psi}_{n}\}_{n=1}^{\infty}$ such that
  1. 1.
    $\|\tilde{\Psi}_{n}\|=1$ for all $n$,
  2. 2.
    $\tilde{\Psi}_{n}$ converges weakly to 0,
  3. 3.
    $\|(\iota U\iota^{-1}-e^{i\lambda})\tilde{\Psi}_{n}\|\to 0$ as $n\to\infty$.
Here, note that the eigenvalue equation $(\iota U_{\pm\infty}\iota^{-1}-e^{i\lambda})\tilde{\Psi}=0$ is equivalent to the transfer matrix recurrence $\tilde{\Psi}(x+1)=T_{\pm\infty}(\lambda)\tilde{\Psi}(x)$.
Also, since the walk has only finitely many defects, $U$ is a finite-rank (hence trace-class) perturbation of the tail operator $U_{\pm\infty}$. Therefore, by the Kato–Rosenblum theorem (see, e.g., [51, 52]), the absolutely continuous spectrum satisfies
$\sigma_{\mathrm{ac}}(U)=\sigma_{\mathrm{ac}}(U_{-\infty})\cup\sigma_{\mathrm{ac}}(U_{+\infty}).$
Theorem 2.
The spectrum of $U$ can be decomposed as
$\sigma(U)=\sigma_{\mathrm{disc}}(U)\,\cup\,\sigma^{\infty}_{\mathrm{ess}}(U)\,\cup\,\sigma_{\mathrm{ac}}(U)$
where:
(i) Discrete spectrum (finite multiplicity).
$\sigma_{\mathrm{disc}}(U)=\sigma_{\mathrm{p}}(U)\setminus\bigl(\sigma_{\mathrm{ess}}^{\infty}(U)\cup\sigma_{\mathrm{ac}}(U)\bigr),$
where $\sigma_{\mathrm{p}}(U)$ is given in Theorem 1, and the multiplicity of each eigenvalue is bounded by $1+N(\lambda)$ (Proposition 2).
(ii) Eigenvalues with infinite multiplicity.
$\sigma^{\infty}_{\mathrm{ess}}(U)=\sigma^{\infty}_{\mathrm{ess}}(U_{+\infty})\cup\sigma^{\infty}_{\mathrm{ess}}(U_{-\infty}),$
where
$\sigma^{\infty}_{\mathrm{ess}}(U_{\pm\infty})=\left\{e^{i\lambda}\ \middle|\ A_{\pm\infty}(\lambda)=0,\ \left(\frac{a^{\pm\infty}_{33}}{a^{\pm\infty}_{11}}\right)^{2}=\frac{a^{\pm\infty}_{12}a^{\pm\infty}_{21}}{a^{\pm\infty}_{32}a^{\pm\infty}_{23}}\right\}.$
(iii) Absolutely continuous spectrum.
$\sigma_{\mathrm{ac}}(U)=\sigma_{\mathrm{ac}}(U_{+\infty})\cup\sigma_{\mathrm{ac}}(U_{-\infty}),$
where
$\sigma_{\mathrm{ac}}(U_{\pm\infty})=\left\{e^{i\lambda}\ \middle|\ A_{\pm\infty}(\lambda)\neq 0,\ |\mathrm{tr}(T_{\pm\infty}(\lambda))|\leq 2\right\}.$
Proof.
The expression of $\sigma^{\infty}_{\mathrm{ess}}(U_{\pm\infty})$ follows directly from Corollary 2 applied to the tail operators $U_{\pm\infty}$. Also, Proposition 2 explicitly gives the expression for $\sigma^{\infty}_{\mathrm{ess}}(U)$ and thus we get $\sigma^{\infty}_{\mathrm{ess}}(U)=\sigma^{\infty}_{\mathrm{ess}}(U_{+\infty})\cup\sigma^{\infty}_{\mathrm{ess}}(U_{-\infty})$.
To determine the absolutely continuous spectrum of the homogeneous tails $U_{\pm\infty}$, we restrict to the non-exceptional regime $A_{\pm\infty}(\lambda)\neq 0$, since $A_{\pm\infty}(\lambda)=0$ only corresponds to eigenvalues in $\sigma^{\infty}_{\mathrm{ess}}(U_{\pm\infty})$. Moreover, if $|\mathrm{tr}(T_{\pm\infty}(\lambda))|>2$, then every solution $\tilde{\Psi}(x)$ grows exponentially on one side, and hence one cannot construct a Weyl sequence. Hence, we assume $|\mathrm{tr}(T_{\pm\infty}(\lambda))|\leq 2$ and $A_{\pm\infty}(\lambda)\neq 0$. Then $T_{\pm\infty}(\lambda)$ has an eigenvalue $\zeta_{\pm\infty}^{\pm}$ with $|\zeta_{\pm\infty}^{\pm}|=1$, so there exists a nontrivial solution $\tilde{\Psi}(x)$, which is uniformly bounded ($\|\tilde{\Psi}(x)\|$ is bounded by some constant for all $x$) of the recurrence $\tilde{\Psi}(x+1)=T_{\pm\infty}(\lambda)\tilde{\Psi}(x)$. Let $\chi_{n}$ be the indicator of the interval $\{-n,\dots,n\}$ and set $\tilde{\Psi}_{n}:=\frac{\chi_{n}\tilde{\Psi}}{\|\chi_{n}\tilde{\Psi}\|}$. Since $\tilde{\Psi}_{n}$ fails to solve the eigenvalue equation only near the two endpoints of the interval, hence $\|(\iota U_{\pm\infty}\iota^{-1}-e^{i\lambda})\tilde{\Psi}_{n}\|=O(n^{-1/2})\to 0$ as $n\to\infty$. Therefore, $\{\iota^{-1}\tilde{\Psi}_{n}\}$ is a Weyl sequence and $e^{i\lambda}\in\sigma_{\mathrm{ac}}(U_{\pm\infty})$.
Finally, since Eq. 10 and Eq. 11, the essential spectrum is given by $\sigma_{\mathrm{ess}}^{\infty}(U)\cup\sigma_{\mathrm{ac}}(U)$ and the discrete spectrum can be characterized as $\sigma_{\mathrm{disc}}(U)=\sigma_{\mathrm{p}}(U)\setminus\bigl(\sigma_{\mathrm{ess}}^{\infty}(U)\cup\sigma_{\mathrm{ac}}(U)\bigr)$ from Eq. 9. ∎
Remark 1.
Note that the above decomposition in Theorem 2 does not rule out eigenvalues embedded in the absolutely continuous band set, i.e. phases $e^{i\lambda}\in\sigma_{\mathrm{p}}(U)\cap\sigma_{\mathrm{ac}}(U)$.
Such embedded eigenvalues can be created by the finite defect region when one tail is propagating at $e^{i\lambda}$ and an additional degeneracy occurs at the defects ($A_{x}(\lambda)=0$ at some defect site), so that an exponentially decaying mode on one side can be matched across the defect region to a propagating mode on the other, producing an $\ell^{2}$ eigenstate at a band energy. Any embedded eigenvalues arising in this way have finite multiplicity. In contrast, if one homogeneous tail has a flat-band eigenvalue at $e^{i\lambda}$ while the other tail is propagating at the same phase, then $e^{i\lambda}$ is an embedded eigenvalue of infinite multiplicity. A complete characterization of the parameter regimes producing embedded eigenvalues is left for future work.
III.3 Example: Homogeneous, one-defect, two-phase model
@@ III.3.1 Homogeneous model
$\mathsf{C}_{x}=\begin{bmatrix}a^{(1,1)}&a^{(1,2)}&a^{(1,3)}\\ a^{(2,1)}&a^{(2,2)}&a^{(2,3)}\\ a^{(3,1)}&a^{(3,2)}&a^{(3,3)}\end{bmatrix}$.
@@
$\sigma_{p}(U)=\begin{cases}\left\{\dfrac{\overline{a^{(3,3)}}}{a^{(1,1)}}\,e^{i\Delta}\right\}&\text{if }\displaystyle\left(\dfrac{a^{(3,3)}}{\overline{a^{(1,1)}}}\right)^{2}=\dfrac{a^{(1,2)}a^{(2,1)}}{\overline{a^{(3,2)}}\,\overline{a^{(2,3)}}},\\[4.30554pt] \emptyset&\text{otherwise,}\end{cases}$
@@
[…]
@@
From the main theorem, we can state that for $\lambda\in[0,2\pi)$, $e^{i\lambda}\neq\frac{\overline{a^{(3,3)}}}{a^{(1,1)}}e^{i\Delta}$, (case $A(\lambda)\neq 0$) is an eigenvalue of $U$ if and only if $|\mathrm{tr}(T(\lambda))|>2$ and there exists $\tilde{\Psi}:\mathbb{Z}\rightarrow\mathbb{C}^{2}\setminus\{\mathbf{0}\}$ such that
@@
for $\tilde{\Psi}$ to be exponentially decaying on both the left and the right. However, since $\zeta^{<}\neq\zeta^{>},$ $\tilde{\Psi}(0)=\mathbf{0}$ and this yields $\tilde{\Psi}=\mathbf{0}$. Thus, only $\frac{\overline{a^{(3,3)}}}{a^{(1,1)}}e^{i\Delta}$ (case $A(\lambda)=0$) could be an eigenvalue. From the main theorem, $e^{i\lambda}=\frac{\overline{a^{(3,3)}}}{a^{(1,1)}}e^{i\Delta}$ is an eigenvalue if and only if $\left(\frac{a^{(3,3)}}{\overline{a^{(1,1)}}}\right)^{2}=\frac{a^{(1,2)}a^{(2,1)}}{\overline{a^{(3,2)}}\overline{a^{(2,3)}}}$ and the corresponding eigenvector becomes $\iota^{-1}\tilde{\Psi}$ such that for each $x\in\mathbb{Z}$, there exists $k_{x}\in\mathbb{Z}$ such that
@@
This result confirms that localization does not generally occur in homogeneous three-state QWs unless specific algebraic conditions are satisfied. Even when an eigenvalue exists, the corresponding eigenstates are fundamentally different from those obtained via the transfer matrix method in the inhomogeneous case: they are not exponentially decaying but instead are compactly supported (nonzero only at finitely many positions). This observation aligns with earlier findings in the literature, such as those derived via Fourier analysis in [29, 53], and further motivates the study of inhomogeneous models, where localization, characterized by exponentially decaying eigenstates, emerges more generically.
@@ III.3.2 One-defect model
The one-defect model introduces a localized perturbation to an otherwise homogeneous QW. Specifically, the coin matrix is modified at a single site, typically taken to be the origin $x=0$, while the coin is kept uniform elsewhere. This minimal breaking of translational invariance is sufficient to drastically alter the spectral and dynamical properties of the system. The one-defect model is particularly important in the study of QWs because it serves as a simple yet rich example of how inhomogeneities lead to localization and bound states. The one-defect model is defined such that $\mathsf{C}_{+\infty}=\mathsf{C}_{-\infty}=\mathsf{C}$, and
$\mathsf{C}_{x}=\begin{cases}\mathsf{C}_{0}&x=0,\\ \mathsf{C}&x\neq 0,\end{cases}$
where we denote each element of $\mathsf{C}$ as
$\mathsf{C}=\begin{bmatrix}a^{(1,1)}&a^{(1,2)}&a^{(1,3)}\\ a^{(2,1)}&a^{(2,2)}&a^{(2,3)}\\ a^{(3,1)}&a^{(3,2)}&a^{(3,3)}\end{bmatrix},\ e^{i\Delta}=\det\mathsf{C}.$
@@
[…]
@@
Figure 2: Probability distribution $\mu_{t}^{(\Psi_{0})}(x)$ for the one-defect Fourier walk with different values of $\theta$ at time 100100 with initial state $\Psi_{0}(0)=[\frac{1}{\sqrt{3}},\frac{i}{\sqrt{3}},\frac{1}{\sqrt{3}}]^{t}$ and $\Psi_{0}(x)=\mathbf{0}$ for $x\neq 0$. All of (a-d) have a peak around the origin, which indicates localization. The result shown in Fig. 1 corresponds to these localization.
@@ III.3.3 Two-phase model
Lastly, we consider the two-phase model, where the time evolution differs between the left and right halves of the lattice. That is, the coin matrix takes two distinct values depending on whether $x<0$ or $x\geqslant 0$. This model has been widely studied in the context of quantum localization and topological insulators, where phase boundaries can give rise to localized edge states. The two-phase model is defined such that $\mathsf{C}_{+\infty}\neq\mathsf{C}_{-\infty}$, and
$\mathsf{C}_{x}=\begin{cases}\mathsf{C}_{+\infty}&x\geqslant 0,\\ \mathsf{C}_{-\infty}&x<0.\end{cases}$
@@
For $\lambda\in[0,2\pi)$, $e^{i\lambda}$ is an eigenvalue of $U$ if and only if there exists $\tilde{\Psi}:\mathbb{Z}\rightarrow\mathbb{C}^{2}$ such that Conditions 1 and 2 are satisfied.
@@
[…]
@@
[…]
@@ IV Numerical analysis on Fourier walk
Figure 4: Probability distribution $\mu_{t}^{(\Psi_{0})}(x)$ for the two-phase Fourier walk with different values of $\theta$ at time 100100 with initial state $\Psi_{0}(0)=[\frac{1}{\sqrt{3}},\frac{i}{\sqrt{3}},\frac{1}{\sqrt{3}}]^{t}$ and $\Psi_{0}(x)=\mathbf{0}$ for $x\neq 0$. All of (a-d) have a peak around the origin, which indicates localization. The result shown in Fig. 3 corresponds to these localization.